MacGyver Corporations information is below: Sales $ 900,000 Profit $36,000 $ 150,000 Contribution margin Investment capital Stockholders equity Plant, property, & equipment $135,000 $ 100,000 $120,000 If the residual income for the year was $9,000, the minimum required rate of return must have been: A. 36% B. 4% C. 15% D. 20%
Answer #1
Minimum required return = Profit-Residual income = 36000-9000 = $27000
Minimum required rate of return =Minimum required return/Investment Capital
Minimum required rate of return = 27000/135000= 20%
Option D 20% is correct